# Diverse Part Synthesis for 3D Shape Creation

### v1

@@ 6 Conclusion, limitations, and future work
Our method has a few limitations arising from our network design. The PCN predicts the position of synthesized parts based on their geometry. Thus, in some cases, e.g., if a part is significantly different from the training samples, the PCN may predict a less than ideal placement. Figure 10 shows a few examples, where the parts with a failure in their placement are in orange. We can see that the chair seat and the airplane tail are placed at slightly inaccurate positions by the PCN. To quantify the occurrence of misaligned parts, we use the PCN to reconstruct the test sets of each category of shapes, each test set comprising $100$ shapes not encountered during training. Following a careful visual inspection, we identify $29$ instances of ShapeNet chairs and $27$ instances of ShapeNet airplanes having notably misplaced parts in their reconstructions, while the reconstructions of the PML chairs do not suffer from this issue. In addition, the quality of the part suggestions depends on the diversity present in the training data. For example, certain types of shapes in the ShapeNet dataset do not posses a great variety of parts, and thus part suggestions for these types of shapes is also limited. In contrast, we obtain a richer variety of part suggestions for shapes in the PML dataset, since the training samples have more variety and gradual differences.
In terms of future work, it may be interesting to replace our part-level AE with an AE that predicts a latent vector for an entire input shape and then predict latent vectors of individual parts from the per-shape latent vector dubrovina2019composite ; hertz2022spaghetti . This would make reconstruction more straightforward and combine the advantages of per-shape and per-part representations into a single method.

### v2

@@ 6 Conclusion, limitations, and future work
Our method has a few limitations arising from our network design. The PCN predicts the position of synthesized parts based on their geometry. Thus, in some cases, e.g., if a part is significantly different from the training samples, the PCN may predict a less than ideal placement. Figure 10 shows a few examples, where the parts with a failure in their placement are in orange. We can see that the chair seat and the airplane tail are placed at slightly inaccurate positions by the PCN. To quantify the occurrence of misaligned parts, we use the PCN to reconstruct the test set of each category of shapes, each test set comprising $100$ shapes not encountered during training. Following a careful visual inspection, we identify $29$ instances of the ShapeNet chairs and $27$ instances of the ShapeNet airplanes having notably misplaced parts in their reconstructions, while the reconstructions of the PML chairs and ShapeNet guitars do not suffer from this issue. In addition, the quality of the part suggestions depends on the diversity present in the training data. For example, certain types of shapes in the ShapeNet dataset do not posses a great variety of parts, and thus part suggestions for these types of shapes are also limited. In contrast, we obtain a richer variety of part suggestions for shapes in the PML dataset, since the training samples have more variety and gradual differences.
In terms of future work, it may be interesting to replace our part-level AE with an AE that predicts a latent vector for an entire input shape and then predict latent vectors of individual parts from the per-shape latent vector [15, 35]. This would make reconstruction more straightforward and combine the advantages of per-shape and per-part representations into a single method. Moreover, due to the design of the PCN, our method is currently limited to shapes decomposed into semantic parts. It would be interesting to adapt our method to datasets with a finer shape decomposition, such as PartNet [63].
Acknowledgments
We thank the anonymous reviewers for their valuable comments and suggestions. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant (2022-04903).